The following reaction was performed in a sealed vessel at 774 ∘C : H2(g)+I2(g)⇌2HI(g) Initially, only H2 and I2 were present at concentrations of [H2]=3.40M and [I2]=2.15M. The equilibrium concentration of I2 is 0.0900 M . What is the equilibrium constant, Kc, for the reaction at this temperature?

I keep getting 14.07 and it says it is incorrect! I am doing (4.12)^2/(1.34)(0.90)
Answer #1

H2(g) + I2(g)⇌ 2HI(g)

inital 3.4 2.15 0

Eqm 3.4-x 2.15-x x

given eqm conc, of I2 = 0.09 M

2.15-x = 0.09 => x= 2.06 M

so eqm conc. are

[H2] = 3.4- 2.06 = 1.34 M

[I2] = 0.09 M

[HI] = 2.06 M

Kc=[HI]^2 / [H2][I2]

= 2.06^2/ 0.06*1.34 = 52.78